An individual spends all his income on two goods X and Y. If with the rise in price of good X, the quantity purchased of good Y remains unchanged, what is the price elasticity of demand for X?
1
Expert's answer
2020-08-19T14:56:45-0400

"PED=(\u2206Q\u00f7\u2206P)\u00d7(P\u00f7Q)"

"\u2206Q=x2-x1"

"\u2206P=P2-P1"


"PED=(x2-x1)\u00f7(P2-P1)\u00d7(P\u00f7Q)"